THEORETICAL LINE STRENGTHS AND TRANSITION MOMENTS IN THE Ã ^{2}B_{1} ← X̃ ^{2}A_{1} ELEC-TRONIC TRANSITION OF NH_{2}

P. R. BUNKER, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada; W. P. KRAEMER, Max Planck Institute for Astrophysics, Postfach 1523, D-85740 Garching, Germany; \underline{PER JENSEN}, FB 9 - Theoretische Chemie, Bergische Universität, D-42097 Wuppertal, Germany.

In a recent publication, Kawakita et al.^{a} call for theoretical calculations of transition moments and line strengths associated with the Ã ^{2}B_{1} ← X̃ ^{2}A_{1} electronic transition of NH_{2}; they need such data in order to improve their analysis of cometary spectra involving this electronic transition. We have used the RENNER program system^{b} to provide the results required. The RENNER calculations are based on ab initio calculations of the potential energy surfaces and the electronic dipole moments and transition moments.^{c} We have optimized the potential energy surfaces in a least-squares fit to experimentally derived term values.

aH. Kawakita, J. Watanabe, D. Kinoshita, S. Abe, R. Furusho, H. Izumiura, K. Yanagisawa, and S. Masuda, Publ. Astron. Soc. Japan 53, L5-L8 (2001).

bSee, for example, P. Jensen, G. Osmann, and P. R. Bunker, in: “Computational Molecular Spectroscopy” (P. Jensen and P. R. Bunker, eds.), Wiley, Chichester, 2000, and references therein.

cP. Jensen, T. E. Odaka, W. P. Kraemer, T. Hirano, and P. R. Bunker, Spectrochim. Acta Part A 58, 763-794 (2002).